OpenStudy (anonymous):

Once your rocket reaches space, the captain adjusts the speed. Your captain goes from 24,000 km/hr to 17,000 km/hr over 90 seconds. What is your acceleration?

OpenStudy (ujjwal):

a= (v-u)/t don't forget to change km/hr to m/s.. (however that also depends upon questions demand or rather say answers demand..)

OpenStudy (stormfire1):

\[V_f -V_i = at\]\[24,000 km/hr = 6.667 x 10^3 m/s\]\[17,000 km/hr = 4.722 x 10^3 m/s\]\[4.722x10^3 m/s - 6.667x10^3 m/s = a (90s)\]\[a = -21.61 m/s^2\]